OpenStudy (anonymous):

0 = 2/3* (x+2)^ (-1/3) how do you solve that?

OpenStudy (anonymous):

\[0=\frac{ 2 }{ 3 }\times(x+2)^{\frac{ -1 }{ 3 }}\]

OpenStudy (anonymous):

Did u mean this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Gimme a sec

OpenStudy (anonymous):

Well there are no sollutions